Regular Meeting 1016
Kokeyama, Sato, Uehara, Ge, Nakano, Nagano, Aso
- 24h Laser Hazard from Oct 15th Monday
Schedule for Phase-2
Report from the Last 2 weeks
- IMC - ASC: PZT mirrors for centering QPDs were installed. PZT efficiencies were measured. PZT1 loop was closed. PZT2 yaw efficiency is 10 times lower.
- MZM - PZTs were replaced for both 1 and 2nd MZIs. Two PZTs for each MZI. Guardian is operational for the 1st MZI. Guardian on the 2nd MZI works, but not fully yet. The OLTs were measured. Both displacements were measured. Calibration on going?
PLS - PMC: temperature loop has gone crazy and MN is working on. Guardian have to have it. The thin film polarizer pair was installed. Power will go up to 5W. RefCav: DC offset cable to AOM was something wrong, to be replaced.
- RAM - no progress
- HPL - Heat sinks with coating arrived and to be installed. HP dump review will be done after this meeting.
- PEM - Fujikawa-kun will hunt the MZM noise.
- IMMT1/IMMT2 - to measure TFs in vac.
- MCI and MCO mirrors will clean with the first contacts before December HP.
- Ge is preparing the visa for the next visit
- Koji Arai will come on the week of 29th.
- Nakano-kun Oct 19th Fri, 22nd Mon
- Kokeyama will work on Detchar in the week of Oct 15th from Mozumi (no mine entrance)
- Kokeyama out for Kashiwa open day on Oct 26th Fri.
- Kokeyama out for Detchar f2f meeting on Oct 29th Mon. Coming back in the afternoon on 30th.
- Kokeyama out for LLO from Nov 6 - Nov 22
- 31st Wednesday